### v1

@@ 4 Speeding up TLWE Encryption via Sample Extraction
4.1 Threat Model and Assumptions
For our threat model, we consider a client or user to be a trusted party capable of performing key generation, encryption, and decryption in trusted environment, and a server to be an untrusted party to which the client would like to offload computation on sensitive data. We discuss the threat protections that our system is able to provide at a high level and our deployment assumptions.
First, we assume that any transfer of secret keys for TFHE can be transferred to the device securely, either before device deployment or over a secure channel. We also assume that the deployed IoT device has access to standard network protection mechanisms for confidentiality, integrity, and freshness of transmitted data. These network security mechanisms are still required on top of homomorphically encrypted ciphertexts since they are necessary for protection against active attacks and secure entire network messages (including any network packet headers) rather than just the ciphertext data itself. In our implementation of TFHE-SBC on Raspberry Pi Zero 2W, for example, communication channels are protected by the TLS protocol.
Furthermore, since no party can learn anything about the underlying data through the encrypted ciphertexts alone, HE is able to provide protection against all aforementioned cloud attackers. HE itself is not robust against malicious cloud attackers, who can manipulate encrypted ciphertexts in the cloud in addition to viewing server state. However, HE can still offer data privacy against these adversaries as well. A remaining threat that we consider is the threat of direct attacks on the device. We consider these threats largely out of scope, since we assume that an attack on the device can easily access the underlying private user data (e.g., sensor samples) or the source of data itself. Since an attacker would already have access to the raw data in this scenario, any additional protections on ciphertexts, keys, or error polynomials would be extraneous. However, if a secret key is stored on the device, a device attacker could theoretically also obtain and use this key to decrypt all previously encrypted information stored in the cloud, thereby expanding the scope of a device attack to a potentially large amount of previously offloaded user data. To mitigate this situation, users can choose to update their device secret keys frequently, thus limiting the interval of data that a device attacker could decrypt using the user’s secret key. This would mean that any data encrypted under the old secret key would no longer be able to be analyzed alongside any data encrypted with the new secret key. However, this is not a problem for private inference applications [8] because they do not store ciphertext in the cloud.
Finally, we note that the post-quantum nature of HE additionally provides some robustness against quantum attacks. In particular, attacks mounted by a quantum-time adversary on HE ciphertexts would not be able to access the underlying private data. For our implementation, we leave quantum attacks out of scope. However, if quantum protection is desired, several existing libraries [26, 27] offer solutions for quantum-secure encryption of messages across an untrusted channel (such as a public network).
4.2 Costs Analysis and Our Proposal
Considering the difference between TLWE and TRLWE configurations, we analyze the computational cost of encryption and the communication cost, i.e., ciphertext size.
Computation costs. TLWE encrypts plaintext per 1-bit and includes $N$ times uniformly distributed random number generation, the inner product of a vector of length $N$, and one time normally distributed random number generation.
TRLWE encryption involves $N$ bit encryption, $N$ times uniformly distributed random number generation, $N$ length polynomial multiplication, and $N$ times normally distributed random number generation.
Therefore, the time complexity of TLWE encryption is $\mathcal{O}(N)$ per bit, and that of TRLWE encryption is $\mathcal{O}(\log N)$ per bit since the time complexity of polynomial multiplication using FFT is $\mathcal{O}(N\log N)$.
Replace TLWE with TRLWE. It is clear that TLWE encryption is computationally more expensive and has a larger ciphertext size than TRLWE encryption. In order to avoid TLWE encryption at the client, we propose to replace TLWE encryption with TRLWE encryption and Sample Extraction [4]. As described in Section 2.2, Sample Extraction converts a single TRLWE ciphertext into $N$ TLWE ciphertexts. In general, Sample Extraction is a component used in the bootstrapping process and is not intended to reduce client-side computational and communication costs. The client sends TRLWE ciphertext to the server, and the server converts it into TLWE ciphertext using Sample Extraction, thereby reducing the client’s execution time and ciphertext size. Since Sample Extraction does not increase the noise, this protocol is also superior to Transciphering [23, 24, 25] that involves the bootstrapping. The protocol of the proposed method is shown below.
  1. 1.
    The trusted party generates the private key, Bootstrapping key, and key switching key, and shares the private key with the client device.
  2. 2.
    The client device encrypts the message with TRLWE and shares the ciphertext with the cloud server. TRLWE encrypts by packing multiple bits together.
  3. 3.
    The cloud server converts TRLWE ciphertext to TLWE ciphertext using Sample Extraction.
  4. 4.
  5. 5.
    The trusted party decrypts the result.
5 Bottleneck Analysis in TRLWE Encryption
In this section, we analyze the bottleneck of TRLWE encryption in TFHEpp, to identify areas for optimization of TRLWE encryption implemented in TFHE-SBC. Among the OpenFHE [28] and TFHEpp [12] libraries that handle TFHE schemes, the TFHEpp library has more functions than the others [29]. Moreover, TFHEpp is considered faster than implementation [30] by the authors of the original TFHE, and our work is the first TFHE framework that enables a client to execute TFHE encryption efficiently on single-board computers. Therefore, we employ the TFHEpp implementation as our baseline.
@@ 5 Bottleneck Analysis in TRLWE Encryption
As described in Section 2.1, TRLWE encryption consists of pseudorandom number generation sampled from the uniform and Gaussian distribution, and polynomial multiplication. Table 4 shows the breakdown of TRLWE encryption on the Raspberry Pi Zero 2W using trlweSymEncrypt function in TFHEpp. “Uniform Sampling” and “Gaussian Sampling” refer to random number generation from their respective distributions, and “PolyMul” refers to polynomial multiplication. The most time-consuming operation is the generation of Gaussian random numbers, which account for 83.1% of the total execution time. Since random numbers sampled from the uniform and Gaussian distribution are generated the same number of times, random numbers sampled from the Gaussian distribution are generated slower than those sampled from the uniform distribution per random number generation. PolyMul has less impact on encryption time than random number generation. We give details about the bottleneck of each component and further discuss their optimization.
5.1 Hardware Dependent Pseudorandom Number Generator
@@
TFHEpp can use /dev/urandom provided by the Linux OS or Randen [^1] https://github.com/google/randen as a cryptographically secure pseudorandom number generator (CSPRNG). Randen is an implementation that requires an AES accelerator, but at this time, AES accelerators are not supported except for Raspberry Pi 5, so Table 4 is the execution time when using /dev/urandom. Random numbers obtained from CSPRNG are used to generate uniform and Gaussian random numbers, so the speed of pseudo-random number generation affects the speed of sampling from uniform and Gaussian distribution. Moreover, it has been reported that random number generation using /dev/urandom slows down the encryption speed [31]. [31] pointed out that TRGSW encryption is 35 times faster on the SBC, which supports AES accelerators, than on the Raspberry Pi 4. It is necessary to implement a CSPRNG that is fast even in environments where AES accelerators are not supported like the Raspberry Pi Zero series.
5.2 Gaussian Random Number Generator
@@
The Gaussian random number generation algorithms include Box-Muller method [32], polar method [33], Kinderman method [34], Monty-Python method [35], and Ziggurat method [36]. In TFHEpp, Gaussian random numbers are generated by the std::normal_distribution function included in the C++ standard library, and the random number generation algorithm is the polar method. The polar method eliminates the trigonometric calculations required by the Box-Muller method but requires resampling with a probability of about 21.5%. This means that the expected number of uniformly distributed random numbers generated for generating a single Gaussian random number is 1.27. The ziggurat method, which has a lower rejection probability than the polar method, can speed up the process.
@@
5.3 Multiplying Polynomials and Applicability of VC4 GPU
For the polynomial multiplication, we enabled the option in TFHEpp to use the implementation for AArch64 of the FFT library SPQLIOS [^2] https://github.com/tfhe/experimental-tfhe/tree/master/circuit-Bootstrapping/src/spqlios. SPQLIOS does not use a real FFT, but rather packs the imaginary part of the input as a complex number, so that a polynomial of length $N$ can be transformed using an FFT with a polynomial of length $\frac{N}{2}$ as input. We observed that this SPQLIOUS optimizes well for CPU-only implementations. Does GPU on Raspberry Pi enable faster polynomial multiplication?
Broadcom VideoCore IV [^3] https://docs.broadcom.com/doc/12358545 is the GPU on the Raspberry Pi. Quad Processing Unit (QPU), the main arithmetic unit of the VC4, has four floating-point arithmetic units and operates as a 16-parallel SIMD over four cycles. The floating-point arithmetic units can simultaneously perform addition and multiplication in a single cycle. Four QPUs make up a slice, and there are three slices in VC4. Each of them operates at 0.25GHz, giving a peak performance of 24GFlops. VC4 uses the host memory and reads and writes by DMA. Of the 512 MB of host memory, VC4 is allocated 128 MB by default. We explore whether the FFT, a component of polynomial multiplication, can be accelerated by VC4.
Table 5: FFT/IFFT Execution and PolyMul Accuracy on VC4
@@ 5.3 Multiplying Polynomials and Applicability of VC4 GPU
The FFT implementation GPU_FFT [37] provided by Raspberry Pi. This provides all power-of-2 FFT lengths between $2^{8}$ and $2^{22}$ points inclusive. GPU_FFT uses a kernel character device known as the “mailbox” for communication between the ARM and the Videocore. The channels include memory allocation, code execution, the getting temperature, firmware revision, etc. For parallelism, VC4 has 12 QPUs; but GPU_FFT only uses 8. It happens that bit-rotated writes can be done very neatly through the Vertex Pipe Memory and DMA controller VDW by 8 QPUs working in parallel.
@@
Accuracy. PolyMul by FFT is known to have a small error due to floating-point arithmetic, but FFT on VC4, which supports only single precision, has additional errors. In the TFHE scheme, since the maximum number of coefficients for PolyMul before rounding to the modulus $q$ is $N*(2^{31}-1)$, the number of bits required for exact PolyMul is $4231+\log_{2}N=$ 42 bits, where $N=2^{10}$. The CPU implementation can retain this maximum value because it can operate in double precision. However, GPU_FFT uses single-precision floats for data and twiddle factors, which does not achieve sufficient PolyMul accuracy. This is because the QPU in VC4 contains two single-precision floating-point units, one for addition and one for multiplication, each operating on a vector of four single-precision numbers. We compare the relative root mean square error (RRMSE) between native PolyMul implementation and PolyMul with FFT / IFFT, defined as $\text{RRMSE}(X,Y)={\sqrt{\frac{1}{n}\sum_{i=1}^{n}|X_{i}-Y_{i}|^{2}}}/{\sqrt{\sum_{i=1}^{n}|X_{i}|^{2}}}$. The native implementation provides exact polynomial multiplication, but its computational cost is $\mathcal{O}(N^{2})$. From Table 5, PolyMul on the single-precision VC4, the RRMSE is closer to 1, which is more inaccurate than PolyMul on CPU. Therefore, we execute PolyMul entirely on the CPU to preserve accuracy, although FFT/IFFT on VC4 is faster.
@@ 6 TFHE-SBC
This section describes TFHE-SBC, which streamlines the client-side operations of the TFHE scheme in SBCs based on the bottleneck analysis in Section 5. TFHE-SBC consists of the device library and adapter library, using an implementation derived in part from the TFHEpp library. The device library contains optimized functions for TRLWE encryption and is installed on the client side. The adapter library converts the data encrypted in the device library into a format that can be handled by the TFHEpp library and is installed on the server side. An overview of the implementation is shown in Figure 2.
Figure 2: Composition of TFHE-SBC: The SBC performs TRLWE encryption, and the server converts the TRLWE ciphertext into TLWE ciphertext. FFT for the secret key is performed only once and reused for subsequent encryption.
@@ 6.1 Device Library
This section presents the structure of the device library, including optimized random number generation and PolyMul.
Pseudo-random number generation with BLAKE2. In Section 5.1, we showed that pseudo-random number generation by /dev/urandom is the bottleneck. In TFHE-SBC, only the random seed is taken from /dev/urandom, and the random seed is expanded by BLAKE2444https://www.blake2.net/. BLAKE2 is a cryptographic hash function faster than MD5, SHA-1, SHA-2, and SHA-3, yet is at least as secure as the latest standard SHA-3. This is also employed in the OpenFHE library.
Gaussian random number with Ziggurat method. In Section 5.2, we showed that the generation of Gaussian random numbers by the polar method is the bottleneck of TRLWE encryption in TFHEpp. In TFHE-SBC, we implement the ziggurat method for Gaussian random number generation, which is faster than the polar method. The ziggurat method [36] is considered the fastest of the five algorithms listed in Section 5.2. Belonging to the class of rejection sampling algorithms, it relies on an underlying source of uniformly distributed random numbers, typically from a pseudo-random number generator, as well as precomputed tables. With a probability of 98.8%, random numbers can be generated with two floating-point multiplications, and the expected number of uniformly distributed random numbers when generating one Gaussian random number is 1.01, which is less than that of the polar. In TFHE-SBC, Gaussian random number generation is based on the boost::random::normal_distribution function in the boost library [^5] https://www.boost.org/ to implement the ziggurat method.
Memory reuse in PolyMul. In TFHEpp implementation of TRLWE encryption shown in Section 5, the execution time for polynomial multiplication was 0.35%, but when the random number generation acceleration described above is applied, the polynomial multiplication accounts for a larger percentage of the execution time. As long as GPU-enabled FFT is not effective, we consider cost reduction of PolyMul on CPU. In TFHEpp, the FFT implementation has already been refined by SPQLIOS, TFHE-SBC further reduces computation costs by reusing the secret key after the Fourier transform. In SBC, when the encryption function is called multiple times, the uniform noise must be resampled and Fourier transformed each time, but the secret key is Fourier transformed only once and then reused. This strategy is also illustrated in Figure 2.
Algorithm 1 TRLWEtoTLWEs a.k.a, Sample Extraction [4]
@@
0:  $(a_{0,0},\ldots,a_{0,N-1},b_{0})\in\text{TLWE}(\Delta m_{0})$, $\ldots$, $(a_{N-1,0},\ldots,a_{N-1,N-1},b_{N-1})\in\text{TLWE}_{(}\Delta m_{N-1})$
@@ 6.2 Adapter Library
We additionally provide an “adapter” server module to convert data encrypted by TFHE-SBC to be compatible with the TFHEpp library. The adapter library converts TRLWE ciphertexts received from the client into TLWE ciphertexts that can be handled by TFHEpp as shown in Algorithm 1. The conversion process is based on SampleExtractIndex by TFHEpp. This function copies a part of the TRLWE ciphertext array to the TLWE ciphertext. SBC inputs the TRLWE ciphertext $\text{TRLWE}(\Delta M)$ where $M=\sum_{j=0}^{N-1}m_{j}X^{j}\in\mathcal{R}_{p}$, then the adapter library outputs TLWE ciphertexts $\text{TLWE}(\Delta m_{0}),\text{TLWE}(\Delta m_{1}),\ldots,\text{TLWE}(\Delta m_{N-1})$. Ciphertext processed by the adapter library can be passed to any TFHEpp-implemented application.

### v2

@@ 5.1 Comparative Analysis of TLWE and TRLWE
5.2 Bottlenecks in TRLWE Implementation
We have demonstrated that TRLWE offers significant advantages over TLWE in terms of both computational efficiency and communication costs. In this section, we analyze the key bottlenecks in TRLWE encryption to inform the implementation in TFHE-SBC framework.
@@ 5.2 Bottlenecks in TRLWE Implementation
Table 4 shows the breakdown of TRLWE encryption on the Raspberry Pi Zero 2W using trlweSymEncrypt function in TFHEpp. “Uniform Sampling” and “Gaussian Sampling” refer to random number generation from their respective distributions, and “PolyMul” refers to polynomial multiplication. The most time-consuming operation is the generation of Gaussian random numbers, which account for 83.1% of the total execution time. Since random numbers sampled from the uniform and Gaussian distribution are generated the same number of times, random numbers sampled from the Gaussian distribution are generated slower than those sampled from the uniform distribution per random number generation. PolyMul has less impact on encryption time than random number generation. We give details about the bottleneck of each component and further discuss their optimization.
@@
Hardware dependent PRNG. TFHEpp can use /dev/urandom provided by the Linux OS or Randen [^1] https://github.com/google/randen as a cryptographically secure pseudorandom number generator (CSPRNG). Randen is an implementation that requires an AES accelerator, but at this time, AES accelerators are not supported except for Raspberry Pi 5, so Table 4 is the execution time when using /dev/urandom. Random numbers obtained from CSPRNG are used to generate uniform and Gaussian random numbers, so the speed of pseudo-random number generation affects the speed of sampling from uniform and Gaussian distribution. Moreover, it has been reported that random number generation using /dev/urandom slows down the encryption speed [31]. [31] pointed out that TRGSW encryption is 35 times faster on the SBC, which supports AES accelerators, than on the Raspberry Pi 4. It is necessary to implement a CSPRNG that is fast even in environments where AES accelerators are not supported like the Raspberry Pi Zero series.
@@
Gaussian random number generator. The Gaussian random number generation algorithms include Box-Muller method [32], polar method [33], Kinderman method [34], Monty-Python method [35], and Ziggurat method [36]. In TFHEpp, Gaussian random numbers are generated by the std::normal_distribution function included in the C++ standard library, and the random number generation algorithm is the polar method. The polar method eliminates the trigonometric calculations required by the Box-Muller method but requires resampling with a probability of about 21.5%. This means that the expected number of uniformly distributed random numbers generated for generating a single Gaussian random number is 1.27. The Ziggurat method, which has a lower rejection probability than the polar method, can speed up the process.
@@
5.3 Applicability of GPUs to Polynomial Multiplication
For the polynomial multiplication, we enabled the option in TFHEpp to use the implementation for AArch64 of the FFT library SPQLIOS [^2] https://github.com/tfhe/experimental-tfhe/tree/master/circuit-Bootstrapping/src/spqlios. SPQLIOS does not use a real FFT, but rather packs the imaginary part of the input as a complex number, so that a polynomial of length $N$ can be transformed using an FFT with a polynomial of length $\frac{N}{2}$ as input. We observed that SPQLIOUS optimizes well for CPU implementations, but we also investigate if the Raspberry Pi’s GPU could enable faster polynomial multiplication.
Broadcom VideoCore IV [^3] https://docs.broadcom.com/doc/12358545 is the GPU on the Raspberry Pi, giving a peak performance of 24GFlops. Quad Processing Unit (QPU), the main arithmetic unit of the VC4, has four floating-point arithmetic units and operates as a 16-parallel SIMD over four cycles. VC4 uses the host memory and reads and writes by DMA. Of the 512 MB of host memory, VC4 is allocated 128 MB by default. We explore whether the FFT, a component of polynomial multiplication, can be accelerated by VC4.
Table 5: FFT/IFFT execution and PolyMul accuracy on VC4.
@@ 5.3 Applicability of GPUs to Polynomial Multiplication
The FFT implementation GPU_FFT [37] provided by Raspberry Pi. This provides all power-of-2 FFT lengths between $2^{8}$ and $2^{22}$ points inclusive. GPU_FFT uses a kernel character device known as the “mailbox” for communication between the ARM and the Videocore.
@@
Accuracy. PolyMul by FFT is known to have a small error due to floating-point arithmetic, but FFT on VC4, which supports only single precision, has additional errors. In the TFHE scheme, since the maximum number of coefficients for PolyMul before rounding to the modulus $q$ is $N*(2^{31}-1)$, the number of bits required for exact PolyMul is $4231+\log_{2}N=$ 42 bits, where $N=2^{10}$. The CPU implementation can retain this maximum value because it can operate in double precision. However, GPU_FFT uses single-precision floats for data and twiddle factors, which does not achieve sufficient PolyMul accuracy. This is because the QPU in VC4 contains two single-precision floating-point units, one for addition and one for multiplication, each operating on a vector of four single-precision numbers. We compare the relative root mean square error (RRMSE) between native PolyMul without FFT / IFFT and PolyMul with FFT / IFFT, defined as $\text{RRMSE}(X,Y)={\sqrt{\frac{1}{n}\sum_{i=1}^{n}|X_{i}-Y_{i}|^{2}}}/{\sqrt{\sum_{i=1}^{n}|X_{i}|^{2}}}$. The native implementation provides exact polynomial multiplication, but its computational cost is $\mathcal{O}(N^{2})$. From Table 5, PolyMul on the single-precision VC4, the RRMSE is closer to 1, which is more inaccurate than PolyMul on CPU. Therefore, we execute PolyMul entirely on the CPU to preserve accuracy, although FFT/IFFT on VC4 is faster.
@@ 6 TFHE-SBC
This section describes TFHE-SBC, a framework optimizing client-side TFHE operations for SBCs based on our bottleneck analysis in Section 5. Our framework comprises two components: a client-side device library with optimized TRLWE encryption functions, and a server-side adapter library that converts encrypted data to TFHEpp-compatible format. Figure 2 presents this architecture.
6.1 TRLWE to TLWE Conversion Protocol
TLWE encryption is significantly more computationally expensive and produces larger ciphertexts compared to TRLWE encryption as shown in Section 5.1. To mitigate these inefficiencies at the client side, we propose to replace TLWE encryption with TRLWE encryption followed by Sample Extraction [4]. As described in Section 2.2, Sample Extraction converts a single TRLWE ciphertext into $N$ TLWE ciphertexts. While Sample Extraction is traditionally a component used in the bootstrapping process, we repurpose it specifically to reduce client-side computational and communication costs. In our protocol, the client transmits only the TRLWE ciphertext to the server, which then performs the conversion to TLWE ciphertext using Sample Extraction. This approach significantly reduces both the client-side execution time and ciphertext size. Since Sample Extraction does not increase the noise, our protocol offers superior performance compared to Transciphering approaches [23, 24, 25] that require bootstrapping operations. Our proposed protocol consists of the following steps:
  1. 1.
    The trusted party generates the private key, Bootstrapping key, and key switching key, then securely shares the private key with the client device.
  2. 2.
    The client device encrypts the message using TRLWE through our device library and shares the ciphertext with the cloud server.
  3. 3.
    The cloud server converts the TRLWE ciphertext to TLWE ciphertext using Sample Extraction via our adapter library.
  4. 4.
  5. 5.
    The trusted party decrypts the result using the private key.
Figure 2: Architecture of TFHE-SBC: System workflow showing SBC-side TRLWE encryption and server-side conversion to TLWE ciphertext. The implementation features accelerated random number generation and single-execution FFT for secret keys.
@@ 6.2 Device Library
The device library is structured to optimize performance on resource-constrained SBCs with particular focus on efficient random number generation and polynomial multiplication operations that address the critical bottlenecks identified in our analysis.
Pseudo-random number generation with BLAKE2. In Section 5.2, we identified that pseudo-random number generation by /dev/urandom is the bottleneck. In TFHE-SBC, we mitigate this issue by obtaining only the random seed from /dev/urandom and expanding it using BLAKE2444https://www.blake2.net/. BLAKE2 is a cryptographic hash function that outperforms MD5, SHA-1, SHA-2, and SHA-3 in terms of speed, while maintaining security levels at least equivalent to the latest standard SHA-3. This approach is also employed in the OpenFHE library for similar efficiency gains.
Gaussian random number with Ziggurat method. In Section 5.2, we demonstrated that the generation of Gaussian random numbers using the polar method creates a bottleneck in TRLWE encryption within TFHEpp. In TFHE-SBC, we implement the Ziggurat method for Gaussian random number generation, which offers superior performance compared to the polar method. The Ziggurat method [36] is considered the most efficient among the five algorithms discussed in Section 5.2. As a rejection sampling algorithm, it relies on an underlying source of uniformly distributed random numbers and precomputed tables. With a probability of 98.8%, random numbers can be generated using only two floating-point multiplications. The expected number of uniformly distributed random numbers required to generate one Gaussian random number is 1.01, which is lower than that of the polar method. In TFHE-SBC, we leverage the boost::random::normal_distribution function from the Boost library [^5] https://www.boost.org/ to implement the Ziggurat method efficiently.
Memory reuse in PolyMul. In the TFHEpp implementation of TRLWE encryption described in Section 5.2, polynomial multiplication accounted for only 0.35% of the execution time. However, after applying our random number generation optimizations, polynomial multiplication represents a larger percentage of the overall execution time. Given that GPU-accelerated FFT proved ineffective for our use case, we focused on reducing the computational cost of polynomial multiplication on the CPU. While TFHEpp already utilizes the refined FFT implementation from SPQLIOS, TFHE-SBC further reduces computation costs by reusing the Fourier-transformed secret key across multiple encryption operations. Specifically, when the encryption function is invoked multiple times on an SBC, the uniform noise must be resampled and Fourier-transformed for each operation, but the secret key is transformed only once and subsequently reused. This optimization strategy is illustrated in Figure 2.
Algorithm 1 Sample Extraction: TRLWEtoTLWEs [4]
@@
0:  $(a_{0,0},\ldots,a_{0,N-1},b_{0})\in\text{TLWE}(\Delta m_{0})$, $\ldots$, $(a_{N-1,0},\ldots,a_{N-1,N-1},b_{N-1})\in\text{TLWE}(\Delta m_{N-1})$
@@ 6.3 Adapter Library
We additionally provide an “adapter” server module to convert data encrypted by TFHE-SBC to be compatible with the TFHEpp library. The adapter library converts TRLWE ciphertexts received from the client into TLWE ciphertexts that can be handled by TFHEpp, as detailed in Algorithm 1. The conversion process leverages the SampleExtractIndex function from TFHEpp, which selectively copies elements from the TRLWE ciphertext array to construct the TLWE ciphertext. When an SBC inputs a TRLWE ciphertext $\text{TRLWE}(\Delta M)$ where $M=\sum_{j=0}^{N-1}m_{j}X^{j}\in\mathcal{R}_{p}$, the adapter library generates the corresponding TLWE ciphertexts $\text{TLWE}(\Delta m_{0}),\text{TLWE}(\Delta m_{1})$, $\ldots,\text{TLWE}(\Delta m_{N-1})$. The resulting ciphertexts can be seamlessly integrated with any application implemented using the TFHEpp library.